OpenStudy (anonymous):

f(x) has a m=-1/x^2 and passes through P (1/5,12) find the curve y=f(x) in the xy plane

OpenStudy (anonymous):

?

ganeshie8 (ganeshie8):

\(\large \frac{dy}{dx} = \frac{-1}{x^2}\)

ganeshie8 (ganeshie8):

take the integral

ganeshie8 (ganeshie8):

\(\large y = \int \frac{-1}{x^2}dx\)

OpenStudy (anonymous):

1/x +C

ganeshie8 (ganeshie8):

So, \(\large y = \frac{1}{x} + C\) is the required curve ?

OpenStudy (anonymous):

1/x + 1/5 ?

OpenStudy (anonymous):

i am not sure, have to use point

ganeshie8 (ganeshie8):

since the curve goes thru the point P(1/5, 12), plugin x = 1/5 and y = 12 in equation

OpenStudy (anonymous):

12=1/(1/5)+C

OpenStudy (anonymous):

c=7

ganeshie8 (ganeshie8):

\(\large y = \frac{1}{x} + C\) plugin, x = 1/5, y = 12 \(\large 12= \frac{1}{1/5} + C\) \(\large 12= 5 + C\) \(\large 7= C\)

ganeshie8 (ganeshie8):

yes, plug c value in ur equation

OpenStudy (anonymous):

so answer is y=1/x +7

ganeshie8 (ganeshie8):

Yes !

OpenStudy (anonymous):

cool thanks

ganeshie8 (ganeshie8):

u wlc !